OpenStudy (anonymous):

Your piggy bank has a total of 41 coins in it; some are dimes and some are nickels. If you have a total of $3.45, how many nickels do you have? Numerical Answers Expected!

OpenStudy (anonymous):

x = # of dimes y = # of nickels x + y = 41 <---------there are a total of 41 coins 0.10x + 0.05y = 3.45 <-------------there is a total of 3.45 Solve for x and y x = 41 - y 0.10(41 - y) + 0.05y = 3.45 4.1 - 0.10y + 0.05y = 3.45 -0.05y = -0.65 y = 13 x + 13 = 41 x = 28 ANSWER: 28 dimes and 13 nickels